OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin r = 4 cos 5θ

OpenStudy (jdoe0001):

in short.... the check for symmetry for the y-axis, x-axis or the pole or y=x line \(\large \begin{array}{rlll} &replace\quad (\theta,r) \\\hline\\ \frac{\pi}{2}line \ (y)&(-\theta,-r)\\ polar-axis\ (x)&(r,-\theta)\\ pole\ line\ (y=x)&(-r,\theta) \end{array} \) if they yield an equivalent equation, after simplification, they it has symmetry at that line

OpenStudy (anonymous):

Thanks I got it! (: I figured it out a little before you replied. But thanks for helping me out! (: